# FGit: Fault-Guided Fine-Tuning for Code Generation

### v2

@@ Abstract
Index Terms:
LLM, Code Generation, Software Engineering
11footnotetext: Corresponding author
@@ I Introduction
Recently, fine-tuning LLMs using synthetic datasets generated by teacher models has emerged as a popular paradigm for improving code generation capabilities [1, 2, 3, 4, 5]. This paradigm uses teacher models to generate high-quality instruction-response pairs and construct a dataset. These datasets are then used to fine-tune student models with standard SFT method, which uses instructions to guide LLMs to generate outputs matching reference responses by minimizing cross-entropy loss uniformly across all tokens.
Although these LLMs fine-tuned with standard SFT achieve impressive performance on code generation benchmarks such as HumanEval [6], they sometimes generate plausible-looking but incorrect code variants [7, 8]. For example, Llama-3.1-70B-Instruct [9], a model fine-tuned from Llama-3.1-70B using standard SFT, is capable of solving 82.3% of the problems in HumanEval. However, our analysis shows that 34.5% of its failed cases are largely correct and require modifications in only three or fewer locations to be fixed. We conservatively classify these specific instances as failures attributable to deviations in error-prone segments, highlighting the prevalence of such errors in code generation. As shown in Figure 1, in a task to calculate the product of signs and the sum of magnitudes, although the LLM successfully implements logic similar to the correct version—by correctly handling negative numbers and summing magnitudes-it fails to account for zero. We refer to such crucial differences between correct implementations and similar incorrect variants as error-sensitive segments. These error-sensitive segments act as critical decision points in code generation, where even slight deviations can determine the correctness of the output.
@@ V Discussion
Compared to Mutation-Based Method
@@
Traditional mutation-based approaches provide a natural baseline for our method, as they are also designed to introduce errors into code. We use universalmutator [50], a state-of-the-art, language-agnostic, rule-based mutation tool, to generate faulty versions of the code in Evol-Instruct to compare with FGit. We train SemCoder-S using these mutated samples, keeping all hyperparameters consistent. As shown in Table VI, FGit outperforms the mutation-based approach. We attribute this to the fact that LLMs, trained on vast code corpora, generate a more diverse and realistic spectrum of plausible errors than a fixed set of mutation rules, leading to more effective learning.
Potential Overfitting of Standard-SFT
@@
We observe that Standard-SFT occasionally lead to performance degradation compared to the base model. For instance, as detailed in Table I, MagiCoderS-DS exhibit a decline in performance across multiple benchmarks after undergoing Standard-SFT. We hypothesize that this can be attributable to potential overfitting, where the model is trained on a dataset whose data distribution is already familiar from its initial fine-tuning phase. To validate this, we fine-tune models on datasets they have not previously been exposed to. Specifically, we fine-tune MagiCoder-DS on PYX and Evol-Instruct, and SemCoder on OSS-Instruct and Evol-Instruct. As shown in Table VII, the models achieve more gains when trained on unseen datasets. For example, SemCoder, when fine-tuned on either OSS-Instruct or Evol-Instruct, outperforms its results when trained on its original dataset, PYX.
@@ VII-A LLMs for Code Generation
As a momentous milestone, Codex [6] boasting a 12-billion-parameter model demonstrates the extraordinary capability to tackle up to 72% of Python programming problems. After that, a new wave of code generation models, such as AlphaCode [53], CodeGen [54], InCoder [55] and StarCoder [56] are proposed and have shown promising results in the code generation task. Building upon these foundations, more code-focused LLMs emerged, such as Magicoder [2], SemCoder [18], WaveCoder [5] and WizardCoder [1]. These specialized LLMs are typically based on general LLMs in solving domain-specific coding tasks through instruction tuning.
@@ VIII Conclusion
Our code is available: https://anonymous.4open.science/r/FAFT-976B.

### v3

@@ Abstract
11footnotetext: Corresponding author.
@@ I Introduction
Recently, fine-tuning LLMs using synthetic datasets generated by teacher models has emerged as a popular paradigm for improving code generation capabilities [1, 2, 3]. This paradigm uses teacher models to generate high-quality instruction-response pairs and construct a dataset. These datasets are then used to fine-tune student models with standard SFT method, which uses instructions to guide LLMs to generate outputs matching reference responses by minimizing cross-entropy loss uniformly across all tokens.
Although these LLMs fine-tuned with standard SFT achieve impressive performance on code generation benchmarks [4, 5, 6, 7], such as HumanEval [8], they sometimes generate plausible-looking but incorrect code variants [9, 10]. For example, Llama-3.1-70B-Instruct [11], a model fine-tuned from Llama-3.1-70B using standard SFT, is capable of solving 82.3% of the problems in HumanEval. However, our analysis shows that 34.5% of its failed cases are largely correct and require modifications in only three or fewer locations to be fixed. We conservatively classify these specific instances as failures attributable to deviations in error-prone segments, highlighting the prevalence of such errors in code generation. As shown in Figure 1, in a task to calculate the product of signs and the sum of magnitudes, although the LLM successfully implements logic similar to the correct version—by correctly handling negative numbers and summing magnitudes-it fails to account for zero. We refer to such crucial differences between correct implementations and similar incorrect variants as error-sensitive segments. These error-sensitive segments act as critical decision points in code generation, where even slight deviations can determine the correctness of the output.
@@ V Discussion
Analysis of Overfitting Dynamics.
Figure 8: Learning curves for SemCoder-S with FGit on Evol-Instruct. (a) training loss, (b) validation loss, and (c) training loss categorized by token type.
To investigate the possibility of potential overfitting, we analyze the learning dynamics. Specifically, we randomly partition Evol-Instruct into a 9:1 training-validation split and reduce the batch size to obtain a finer-grained view of the loss curves. Figure 8(a) and Figure 8(b) illustrate the learning curves for the SemCoder-S model. Notably, the validation loss for FGit shows a consistent decrease, suggesting FGit mitigates the overfitting issue. We hypothesise that by assigning higher weights to more challenging, error-sensitive tokens, FGit guides the model to focus on harder-to-learn patterns, thus mitigating overfitting. Furthermore, to ensure comprehensive learning across the entire code structure, we assign a non-zero base weight ($\alpha-1$) to non-sensitive tokens. Figure 8(c) confirms that the loss for both sensitive and non-sensitive tokens decreases during training.
Comparison with Mutation-Based Method.
@@
Traditional mutation-based approaches provide a natural baseline for our method, as they are also designed to introduce errors into code. We replace the LLM-based method with the rule-based universalmutator [52] to generate faulty versions of the code in Evol-Instruct. It is a state-of-the-art, language-agnostic mutation tool. We train SemCoder-S using these mutated samples, keeping all hyperparameters consistent. As shown in Table VI, FGit outperforms the mutation-based approach. We attribute this to the fact that LLMs, trained on vast code corpora, generate a more diverse and realistic spectrum of plausible errors than a fixed set of mutation rules, leading to more effective learning.
@@
Potential Overfitting of Standard-SFT. We observe that Standard-SFT occasionally lead to performance degradation compared to the base model. For instance, as detailed in Table III, SemCoder exhibit a decline in performance across multiple benchmarks after undergoing Standard-SFT. We hypothesize that this can be attributable to potential overfitting. The model is trained on PYX whose data distribution is already familiar from its initial fine-tuning phase. To validate this, we fine-tune models on datasets they have not previously been exposed to. Specifically, we fine-tune MagiCoder-DS on PYX and Evol-Instruct, and SemCoder on OSS-Instruct and Evol-Instruct. As shown in Table VII, the models achieve more gains when trained on unseen datasets. For example, SemCoder, when fine-tuned on either OSS-Instruct or Evol-Instruct, outperforms the version trained on its original dataset, PYX.
@@ VII-A LLMs for Code Generation
LLMs are increasingly being leveraged to automate various tasks in software engineering  [55, 56, 57, 58, 59]. Among these, code generation has emerged as a particularly prominent area of research and application [60, 58, 61, 62]. As a momentous milestone, Codex [8] boasting a 12-billion-parameter model demonstrates the extraordinary capability to tackle up to 72% of Python programming problems. After that, a new wave of code generation models, such as AlphaCode [60], CodeGen [58], InCoder [61] and StarCoder [62] are proposed and have shown promising results in the code generation task. Building upon these foundations, more code-focused LLMs emerged, such as Magicoder [7], SemCoder [20], WaveCoder [3] and WizardCoder [1]. These specialized LLMs are typically based on general LLMs in solving domain-specific coding tasks through instruction tuning.
@@ VIII Conclusion
Acknowledgment
This research/project was supported by the National Natural Science Foundation of China (No.62202420 and No. 62302430), Zhejiang Provincial Natural Science Foundation of China (No.LZ25F020003 and No.LQ24F020017).
@@ Acknowledgment
Our code is available: https://github.com/ZJU-CTAG/FGIT.